Suppose that the annual income of a fast-food chain is normally distributed with a mean of $72,000 and a standard deviation of $9,000.

a. What is the probability that a randomly selected fast-food restaurant will generate an annual income between $65,000 and $75,000?

b. What is the probability that a randomly selected fast-food restaurant will generate an annual profit of more than $80,000?

c. What minimum income does a fast-fast restaurant need to earn to be in the top 5% of incomes?

d. What maximum income does a fast-food restaurant need to earn to be in the bottom 30% of incomes?


1
Expert's answer
2021-04-29T18:04:40-0400

mean "= \mu = 72000"

standard deviation "= \sigma = 9000"

a.

"P(65000<X<75000) = P(\frac{65000-7200}{9000}<Z< \frac{75000-72000}{9000}) \\

= P(-0.777<Z<0.333) \\

= P(Z<0.333) -P(Z< -0.777) \\

= 0.6304 -0.2185 \\

= 0.4119"

b.

"P(X>80000) = 1 -P(X<80000) \\

= 1 - P(Z< \frac{80000-72000}{9000}) \\

= 1 -P(Z<0.888) \\

= 1 -0.8127 \\

= 0.1873"

c.

"P(Z> \frac{x-72000}{9000}) = 0.05 \\

P(Z>1.645) = 0.05 \\

\frac{x-72000}{9000} = 1.645 \\

x -72000 = 14805 \\

x = 86805"

d.

"P(Z< \frac{x-72000}{9000}) = 0.3 \\

P(Z< -0.524) = 0.3 \\

\frac{x-72000}{9000} = -0.524 \\

x -72000 = -4716 \\

x = 67284"